OpenStudy (anonymous):

Write and simplify the equation you would use to solve the problem below. You buy 5 tickets to a concert for you and your friends. You pay $170, including a $20 service charge. How much does one ticket cost? A. 5*20+x= 170 B. 20x + 5 = 170 C. 5(20 + x) = 170 D. 5x + 20 = 170

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

You pay a $20 service charge: 20 You buy 5 tickets, at $x/ticket: 5x No other charges, so add them up and set equal to the amount paid:

OpenStudy (anonymous):

B?

OpenStudy (whpalmer4):

B contains "20x + 5 = 170" I gave you "20" and "5x"

OpenStudy (whpalmer4):

add up "20" and "5x" and set equal to 170

OpenStudy (anonymous):

20 20+5x= 170?

OpenStudy (anonymous):

*So

OpenStudy (whpalmer4):

sorry, mistyped that! yes, \[20+5x= 170\] it may appear in a different order in the answer because a + b = b + a...

OpenStudy (anonymous):

Okay. (:

OpenStudy (anonymous):

Thanks. (:

OpenStudy (whpalmer4):

so, what's your answer, and how much does 1 ticket cost?

OpenStudy (anonymous):

its not B